OpenStudy (anonymous):

How do you solve Find two consecutive integers such that four times the greater added to the lesser is 98

OpenStudy (anonymous):

There is no solution. \[n+4(n+1)=98\\ 5n+4=98\\ 5n=94\\ n=\frac{94}5\]As you can see, this is not an integer.

OpenStudy (anonymous):

So what would you have to do after that

OpenStudy (anonymous):

There exist no integers that fit the constraint!